OpenStudy (anonymous):

What Is The Ratio Of X To Y? 2/3= x ______ 3x+y

OpenStudy (cggurumanjunath):

cross multiply !

OpenStudy (cggurumanjunath):

2(3x+y) =3x

OpenStudy (cggurumanjunath):

6x+2y = 3x

OpenStudy (cggurumanjunath):

6x-3x=-2y

OpenStudy (cggurumanjunath):

3x=-2y

OpenStudy (cggurumanjunath):

x/y = -2/3